lisa is six years older than susan,jessie is 5 years older than lisa ,and the total of their ages is 41 how old is susan
Answer by ankor@dixie-net.com(22740)

let the 3 girls be, L, S, & J
write an equation for each statment
" lisa is six years older than susan,"
L = S + 6
or
S = L - 6
:
"jessie is 5 years older than lisa ,"
J = L + 5
:
" the total of their ages is 41"
L + S + J = 41
replace S with (L-6); replace J with (L+5)
L + (L-6) + (J+5) = 41
3L - 1 = 41
3L = 42
L = 42/3
L = 14 yrs is Lisa's age
:
how old is susan
S = 14 - 6
S = 8 yrs is Susan